40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3699     Entry time: Tue Oct 12 17:42:57 2010
Author: yuta 
Type: Update 
Category: SUS 
Subject: very first measurement of Q-values for MC1 

Background:
 Data aquisition system is fixed, and now we can use the Dataviewer to measure Q-values of the ringdowns for each DOF, each optics.
 First of all, I measured MC1 suspention damping servo for a test.

What I did:
1. Used DAQ channels activated in this entry(#3690) to see and compare the ringdowns when the damping servo is on and off with the Dataviewer.

2. Plotted the data and fitted the ringdown using this formula;
  p[0]*exp(-p[1]*t)*sin(p[2]*t+p[3])+p[4]
 I used python's scipy.optimize.leastsq for the fitting.

3. Calculated the resonant frequency f0 and Q-value using following formulas;
  f0=2*pi*sqrt(p[1]**2+p[2]**2)
  Q=f0/(2*pi)/(2*p[1])

4. For plotting, I subtracted the offset(=p[4]).

All parameters I used for this measurement are automatically saved here;
  /cvs/cds/caltech/burt/autoburt/snapshots/2010/Oct/12/13:07/c1mcs.epics
(-1,0,1 for all matrix elements, GAIN=3,3,3,150 for POS,PIT,YAW,SIDE)

Result:
 Attached is the plot of each 4 DOF ringdown when servo is off and on.
 "servo off" means off for that DOF. Servo for the other 3 DOFs are on.

 As you can see clearly, the damping servo is working.

 The resonant frequencies and Q-values calculated from the fitting are as follows;

  servo off servo on
f0 (Hz) Q f0 (Hz) Q
POS 0.97 large 0.97 16
PIT 0.71 96 0.73 6.9
YAW 0.80 100 0.82 8.9
SIDE 0.99 large 0.99 27


 Resonant frequencies and Q-values have about 1% and 10% error respectively.
 I estimated it from my 2-time measurement of the POS ringdown.

Next work:
 - Find and modify some scripts to optimize the matrix elements
 - Calibrate the displacement
 - Do the same thing for other optics

Attachment 1: MC1ringdown.png  473 kB  | Hide | Hide all
MC1ringdown.png
ELOG V3.1.3-